OpenStudy (anonymous):

Help please??? Write the equation of an ellipse with vertices (10, 0) and (-10, 0) and co-vertices (0, 2) and (0, -2).

OpenStudy (anonymous):

\[\frac{ x^2 }{ \sqrt{10} }+\frac{ y^2 }{ \sqrt{2} }\]

OpenStudy (anonymous):

\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }\] is the standard form of an ellipse. After plugging in a (+/- 10) and b (+/- 2) you simplify which is how you get your answer

OpenStudy (anonymous):

so is your first reply the answer? @ellecim_m

OpenStudy (anonymous):

yes